Canyouhelpme with this questionYou are managing a project to expand a major county road in your area. The project is expected to take 1 year and the expenses are budgeted to be $12,500 per month. Six months after the project started, you find out that the project is 50% complete, and that you have spent $70,000. Fill in the information in the table below:

Item Value
BAC
PV
EV
AC
CV = EV-AC
SV = EV – PV
CPI = EV/AC
SPI = EV/PV
EAC = BAC/CPI
ETC = EAC-AC
VAC = BAC-EAC
TCPI (Cost) = (BAC-EV)/(BAC-AC)

Based on the values you computed, is the project ahead of schedule or behind schedule? Will the project be completed under budget or will it exceed the budget?
Answer #1

BAC = 12500 * 12 = 150,000
PV = 50% of BAC = 75,000
EV = 50% of BAC = 75,000
AC = 70,000
CV = EV-AC = 5,000
SV = EV-PV = 0
CPI = EV/AC = 1.071
SPI = EV/PV = 1
EAC = BAC/CPI = 140,000
ETC = EAC-AC = 70,000
VAC = BAC-EAC = 10,000
TCPI = (BAC-EV)/(BAC-AC) = 0.9375

NOTE:

BAC:
BAC is the estimated value of what the project will cost.

PV:
The formula to calculate Planned Value is simple. Take the planned percentage of the completed work and multiply it by the project budget and you will get Planned Value.

Planned Value = (Planned % Complete) X (BAC)

AC:
This is the second element of earned value management. Actual Cost is the total cost incurred for the actual work completed to date. Simply put, it is the amount of money you have spent to date.

EV:
As per the PMBOK Guide, “Earned Value (EV) is the value of work performed expressed in terms of the approved budget assigned to that work for an activity or WBS component.”
Earned Value = % of completed work X BAC

I hope this helped you. Let me know if you need any assistance or help. I shall be glad to help you.